OpenStudy (anonymous):

Find RT if RT=x-6and the measure of angle PTR is 8x-6

OpenStudy (anonymous):



OpenStudy (mathlegend):

What chapter is this in Geometry or what section?

OpenStudy (anonymous):

I don't know because this is a review problem... I took geometry years ago... but the worksheet is titled Bisectors, Medians, and Altitudes

OpenStudy (anonymous):

Any thoughts?

OpenStudy (mathlegend):

No, I'm trying to think though.

OpenStudy (anonymous):

OK

OpenStudy (mathlegend):

I think I got it...

OpenStudy (mathlegend):

PTR is a right triangle... this is dealing with altitudes. We know this is a right triangle because PT is the altitude.

OpenStudy (mathlegend):

PTR = 90

OpenStudy (anonymous):

Oh...I think I get it now, thank you so much!

OpenStudy (mathlegend):

8x - 6 = 90

OpenStudy (mathlegend):

8x-6=90 +6 +6 ----------- 8x = 96 x = 12

OpenStudy (mathlegend):

RT = x-6 12-6 RT = 6

OpenStudy (anonymous):

OK thanks!

OpenStudy (mathlegend):

I had read up on this from my downloaded Geometry textbook. :)